OpenStudy (sparklestaraa):

Choose the correct description of the graph of the compound inequality: 3x + 1 > 7 and 4x >= 24 (4 A number line with an open circle on 2, a closed circle on 6, and shading in between A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right A number line with a closed circle on 2, an open circle on 6, and shading in between

OpenStudy (sparklestaraa):

I chose A is it right @dtan5457

Nnesha (nnesha):

can you check your question if that sign are right ???

Nnesha (nnesha):

is that \[4x \ge 24 ????? or \le \]

Nnesha (nnesha):

or is that 3x + 1 > 7 ?? or < check that please

OpenStudy (dtan5457):

4x >= 24 (4 ?

OpenStudy (dtan5457):

rephrase please

OpenStudy (dtan5457):

is it just 4x >= 24?

OpenStudy (dtan5457):

\[3x+1> 7\] \[4x \ge 24\]

OpenStudy (dtan5457):

are these correct?

OpenStudy (sparklestaraa):

no 4x <= 24

OpenStudy (sparklestaraa):

1st is right :D

Nnesha (nnesha):

see

OpenStudy (sparklestaraa):

@dtan5457

Nnesha (nnesha):

so yes then your answer is right

Nnesha (nnesha):